OpenStudy (anonymous):

Can someone check my work?? (2x-7)^2\ I got: -24x+49

OpenStudy (anonymous):

Ignore the Slash

OpenStudy (luigi0210):

how did you get the -24?

OpenStudy (anonymous):

I foiled

OpenStudy (johnweldon1993):

 That's 'foiling' right? try it one more time...

OpenStudy (luigi0210):

you also don't have a squared term..

OpenStudy (johnweldon1993):

so we do 2x * 2x... =? 2x * -7... =? -7 * 2x... =? -7 * -7... =?

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

and -14 for the middle two

OpenStudy (johnweldon1993):

When you multiply 2 'x' numbers....the x becomes squared.... 2x * 2x = 4x²

OpenStudy (anonymous):

Oh yes right

OpenStudy (johnweldon1993):

and when you multiply a constant times an 'x' number...you multiply the numbers...but keep the x 2x * -7 = -14x right?

OpenStudy (anonymous):

yeah

OpenStudy (johnweldon1993):

and then just the -7 times -7 is?

OpenStudy (anonymous):

49

OpenStudy (johnweldon1993):

Right ...so altogether you have 4x² - 14x - 14x + 49 or 4x² - 28x + 49 right?

OpenStudy (anonymous):

ok I see what I did wrong

OpenStudy (johnweldon1993):

There you go :)

OpenStudy (anonymous):

Thank you :)

OpenStudy (johnweldon1993):

anytime!